OpenStudy (naga447):

Can someone please check my work? 1. Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (naga447):

(0.5+0.2)x – 0.5x = 500 0.2 is the profit of reselling the flowers so each flower costs 0.7 x = the number of flowers that they bought and sold (0.5+0.2)x – 0.5x = 500 0.7x – 0.5x = 500 0.2x = 500 0.2 * 2500 = 500 x = 2500

OpenStudy (anonymous):

yes,you're correct :)

OpenStudy (naga447):

@PFEH.1999 Thanks!